# Genixer: Empowering Multimodal Large Language Model as a Powerful Data Generator

### v2

@@ Abstract
Instruction tuning data is essential for training the Multimodal Large Language Models (MLLMs). However, the creation of high-quality instruction tuning data presents significant challenges. Prior methods that depended on GPT-4 for data generation were not only costly but also lacked satisfactory performance in complex tasks (i.e., grounding-based reasoning tasks). To address these issues, we developed an innovative data generation pipeline Genixer to generate various high-quality instruction tuning data, including nine representative tasks, e.g., Common VQA, REC, REG, and PointQ. Specifically, Genixer provides a unified solution with four key steps for alleviating the difficulty of data generation: (i) instruction data collection, (ii) instruction template design, (iii) empowering MLLM, and (iv) data generation and filtering. Subsequently, the superior qualitative results of our Genixer demonstrate that current MLLMs have a strong potential to evolve into powerful data generators. Additionally, to validate the efficacy of generated data quantitatively, we add the instruction tuning data produced by Genixer into the training of two representative MLLMs and observe the consistent improvements on various VQA tasks and multimodal benchmarks. The data, code, and models will be released.
@@
Multimodal Large Language ModelInstruction Tuing1 Introduction
Large Language Models (LLMs) [6, 37, 2, 14] have achieved remarkable success in understanding human instructions and demonstrated their high proficiency in tackling complex natural language tasks. This progress has recently extended to Multimodal Large Language Models (MLLMs) [17, 40, 13, 29, 53, 14, 23, 9, 51, 45, 55, 3], where visual information is encoded by a dedicated visual encoder and seamlessly integrated into LLMs. This integration empowers MLLMs with the ability to analyze and reason about multimodal tasks effectively.
The training of a satisfactory MLLM heavily relies on the availability of sufficient high-quality visual instruction tuning data. The data are, in many ways, the lifeblood of contemporary general-purpose intelligence systems, much like oil powers industrial processes. To attain superior performance, one pioneer model like InstructBLIP [13] requires training on multiple held-in datasets, such as VQAv2 [15] and OKVQA [35], from traditional Vision-Language (VL) tasks. Unfortunately, these held-in datasets suffer from a limitation in image diversity, as most of them originate from the COCO dataset [27], potentially restricting the model generalization ability on unseen visual data. This naturally raises the fundamental question: how can we obtain a sufficient quantity of high-quality instruction tuning data for training MLLMs?
As of now, there is no effective solution to this challenge yet. The traditional approach of manual labeling proves prohibitively expensive and impractical, particularly since MLLM training often necessitates million-scale instruction tuning data to ensure robust generalization. Another potential solution is to leverage powerful models to generate high-quality data. Unfortunately, almost all open-sourced pretrained MLLMs fall short in this regard. As exemplified in Fig. 1, models like LLaVA-1.5 [28], mPLUG-owl [51], and Shikra [9] tend to treat conversation generation as a dense captioning task and struggle to produce natural and coherent dialog data and IntructBLIP [13] directly fails. Contrarily, our produced dialogue exhibits both versatility and intelligence, attaining a level comparable to that of GPT-4V [38]. Except for leveraging current MLLMs, another viable way is to utilize GPT-4V with carefully designed prompts via the released API, but such a choice leads to high financial costs. For instance, producing millions of samples, each with thousands of tokens, with GPT-4V [38] would cost millions. Moreover, as illustrated in Fig. 2, GPT-4V cannot generate satisfied instruction tuning data of some complex data types (e.g., referring expression comprehension), further limiting its usage for exclusive model learning.
In this work, we address the challenge of generating highly qualified multimodal data. To this end, we introduce Genixer, an automatic data generation pipeline that provides a unified solution with four key steps: (i) instruction data collection, (ii) instruction template design, (iii) empowering MLLM, and (iv) data generation and filtering. We begin by categorizing current Vision-Language (VL) tasks into general and grounding tasks. These categories include nine data types, e.g., Common VQA, Multi-choice VQA, Referring Expression Comprehension (REC), and Referential Dialogue (RD). See details in Sec. 3.1. Then, we compile datasets corresponding to these nine tasks and consistently reformat the input data into the instruction-following format. To acquire the flexibility of automatic large-scale and user-guided data generation, we propose a two-level instruction mechanism for creating data in two modes: type-agnostic generation and type-specific generation (see Sec. 3.2). In the type-agnostic mode, given one image, users can pose an instruction without specifying a particular data type to allow the model to perform data generation freely. This is because, in some cases, preemptively determining the optimal data type for generation becomes particularly daunting, especially when it involves large-scale data production for model training purposes. For the type-specific mode, users can ask the model to generate a specific data type from the above nine data types. Subsequently, we leverage two representative models, LLaVA1.5 [28] and Shikra [9] evolve them into the intelligent data generators with the abovementioned two-level instruction template yielding two data generators, Genixer and Genixer, for general and grounding data generation, respectively (See Sec. 3.3). To maintain high data quality, we employ the third-party MLLM Fuyu-8B [5] as an evaluator within our specially designed evaluation framework for general task evaluation, detailed in Sec. 3.4. Additionally, we utilize CLIP [19] to calculate the region-expression similarity for REC-like data, enabling us to filter out samples exhibiting low similarity.
Experimental results show that using 915K VQA-like tuning data generated by Genixer to train LLaVA1.5 can make a significant performance enhancement compared to the vanilla LLaVA1.5 across a diverse set of tasks, such as 3.8% improvement on Vizwiz [16] and 2.7% on SicenceQA [31]. Furthermore, our Genixer extends its utility to generate instruction-tuning data for visual grounding tasks akin to those encountered in the SoTA grounding model, Shikra. These results underscore the remarkable efficacy of our proposed data generation pipeline, Genixer. In summary, our contributions are:
We introduce an innovative multimodal data generation pipeline, Genixer, that enables existing MLLM models to function as potent data generators to effectively and affordably generate high-quality instruction data.
We contribute two data generators Genixer and Genixer for a wide range of multimodal instruction tuning data generation.
We contribute two high-quality multimodal data: 915K VQA-like data, which can greatly improve LLaVA1.5 on various tasks like Vizwiz and ScienceQA, and 350K REC-like data that enhance Shikra on the grounding tasks like REC datasets.
@@ 2 Related Work
Multimodal Large Language Models. Large Language Models (LLMs) showcased the remarkable complicated reasoning abilities. Some well-known open-souced LLMs include FlanT5 [12], OPT [54], LLaMA [46], Vicuna [11] and LLaMA-2 [47] show exceptional reasoning ability of solving math, codes problems. By leveraging these LLMs, a surge of multimodal modes [25, 29, 13, 3, 9, 49, 55, 8, 7, 32, 50, 17, 40, 36, 51, 14] are proposed to integrate the visual information for diverse multimodal reasoning tasks such as image captioning [10, 1, 52] and visual question answering [15, 35, 18, 16]. LLaVA [29] is a pioneering approach that adopts a single linear layer to project the CLIP [42] encoder extracted visual features to the input of LLM. Different from LLaVA, InstructBLIP [13] employs an instruction-aware feature extractor and obtains advanced performance on various tasks building upon the pretrained BLIP2 [25]. Besides focusing on traditional multimodal tasks, Shikra [9] and PVIT [7] pay attention to the grounding multimodal tasks by carefully designing the visual instruction data or employing the region-based vision encoder. Except for these MLLMs, CogVLM [49], Qwen-VL [3], and Kosmos-2 [40] explore adopting a billion-scale pertaining data corpus to enhance the model generalization ability and robustness.
Multimodal Instruction Data. High-quality multimodal instruction data is crucial for training a robust MLLM. Two main setups are as follows: 1) Transforming the current vision-language datasets into the instruction-tuning format, e.g., InstructBLIP. Such choice is limited by the diversity of image sources. 2) Some approaches LLaVA, VisionLLM[50], and Shikra, resort to prompting the GPT-4 [37] language model to generate corresponding instruction data. This way requires the image datasets to have enough captions or region-based annotations (e.g., bounding boxes), which heavily restricts the data scale. Additionally, prompting commercial LLMs incurs high costs, and even GPT-4V [38] may not address the data generation effectively on some specific tasks, as illustrated in Fig. 2. To this end, we introduce Genixer, an innovative pipeline that explores the capabilities of MLLMs to generate high-quality multimodal instruction data.
3 Genixer: An Automatic Instruction Tuning Data Generation Pipeline
@@
General | Common VQA | VQAv2, GQA, Counting110K, POPE | 583K |
Adv VQA | POPE | 20K | |
@@
Though current MLLMs show exceptional capability of handling various multimodal tasks [28, 3, 24], rare works are concentrating on exploring the data generation capability. In this work, instead of focusing on how to make an MLLM a good problem-solver, we try to provoke the potential ability of current MLLM to generate instruction tuning data automatically and intelligently. To this end, We propose Genixer, as illustrated in Fig. 3, which is a novel pipeline that contains four key steps, including 1) instruction data collection, 2) instruction template design, 3) empowering MLLMs, and 4) data generation and filtering. In the following, we will elaborate on these four key steps.
@@ 3.1 Instruction Data Collection
In accordance with the prevalence and practical relevance of real-world multimodal tasks, we carefully select nine representative multimodal tasks as listed in Tab. 1 for training Genixer, including Common Visual Question Answering (Common VQA), Adversarial-based VQA (Adv VQA), Multi-choice VQA (MC VQA), Multi-turn Dialogue (MD), Referring Expression Comprehension (REC), Referring Expression Generation (REG), PointQA, QCA and Referential Dialogue (RD). We divide these tasks into two categories, general tasks and grounding tasks, and display some examples in Fig. 4 for visual inspection.
@@ 3.2 Instruction Template Design
In an automatic data generation context, where image content is agnostic, preemptively determining the optimal data type for generation becomes particularly daunting, especially when it involves large-scale data production for model training purposes. For instance, in a landscape picture without any humans or objects, it is hard to ask a high-quality multi-choice question, and it is easy to generate an adversarial question such as “Is there a boat in the image?” Hence, it is imperative to delineate two key modes for a robust data-generative MLLM: 1) type-agnostic data generation and 2) type-specific data generation.
Two-level Instructions. To achieve this goal, we propose a two-level instruction template for achieving controllable question-answer generation at the inference stage. Specifically, the complete instruction design is as follows:
@@
Regarding General Instruction, we meticulously compile a list of instructions, and for each piece of training data, we randomly choose one example to use during the training process. For instance, “Please provide a clear and direct question and answer after examining the image”. This instruction plays a role in allowing the model to generate diverse types of instruction tuning data without any specific condition referred to as mode 1.
Then, for Specific Instruction, it denotes designating the specific question type, such as “This is a Common VQA task”, allowing us to control the output type of the generated data referred to as mode 2.
Controlling Constant. During the training phase, we set a constant for controlling the ratio of training samples that are exclusive with General Instruction. Consequently, in the inference phase, the model is able to switch the mode by adding specific instructions or not. For example, as illustrated in Fig. 5, the model is capable of generating various types of data in the absence of specific instructions. Simultaneously, it can produce specific outputs when guided by a detailed prompt, like “This is an MC VQA task”.
@@ 3.3 Empowering Current MLLMs
A robust model with multimodal reasoning ability is the cornerstone for building an instruction tuning data generator. Therefore, we resort to delineating the current remarkable MLLM as the base model and then fine-tuning it with our developed two-level instruction template. Consequently, we present two data generators, Genixer and Genixer, to respectively generate instruction tuning data for the two data categories, general tasks, and grounding tasks as summarized in Tab. 1.
Overall Framework of Genixer and Genixer. For brevity, we denote the MLLM model as , the input general and specific instructions as and , respectively. Then, given an image , our target is to enforce to generate question and corresponding answer :
@@
where is the whole sentence compose and , and is current prediction token. is the length of the model response sequence. denotes the total trainable parameters in (e.g., the parameters of projector and LLM with LLaVA1.5 backbone). In the following, we select LLaVA1.5 and Shikra to generate two kinds of data because of their SoTA performance in the corresponding areas. It is noteworthy that Genixer is flexible and applicable to work with any strong MLLM.
Training of Genixer. Genixer trains the pretrained LLaVA1.5 [28], a SoTA model for four kinds of general tasks, including Common VQA, Adv VQA, MC VQA, and MD. As summarized in Tab. 1, we only sample a subset of data for these four data types to improve training efficiency. The set to 0.2, 0.2, 0.5, 0.2, respectively. The different ratios are because of the different data sizes of these four data types. To keep the training data balanced, we handcrafted the values. Finally, we use AdamW [30] optimizer with a learning rate of and a batch size of 128 to fine-tune the pretrained LLaVA1.5 for one epoch which takes about 14 hours.
Training of Genixer. Here, we aim to train a data generator for grounding tasks such as region-based VQA that involve the object coordinates by referring to the corresponding objects. To ensure the generation, we utilize the entirety of the RefCOCO [20] and VG [21] datasets. Nonetheless, as indicated in Tab. 1, the dataset size for RD is relatively small. To counteract potential biases in the generation, Genixer that finetunes a SoTA model Shikra [9] on region-based tasks adopts two-phased training. The first phase focuses on the REC and REG data generation. In the second phase, Genixer adds PointQA, QCA, and RD data while deliberately reducing REC and REG data for data trade-off. To train the Genixer, we set the equal to 0.5 for each data type. We utilize the AdamW optimizer, applying learning rates of for the first phase and for the second phase. The batch sizes are set to 128 and 64, respectively.

### v6

@@ Abstract
Multimodal Large Language Models (MLLMs) demonstrate exceptional problem-solving capabilities, but few research studies aim to gauge the ability to generate visual instruction tuning data. This paper proposes to explore the potential of empowering MLLMs to generate data independently without relying on GPT-4. We introduce Genixer, a comprehensive data generation pipeline consisting of four key steps: (i) instruction data collection, (ii) instruction template design, (iii) empowering MLLMs, and (iv) data generation and filtering. Additionally, we outline two modes of data generation: task-agnostic and task-specific, enabling controllable output. We demonstrate that a synthetic VQA-like dataset trained with LLaVA1.5 enhances performance on 10 out of 12 multimodal benchmarks. Additionally, the grounding MLLM Shikra, when trained with a REC-like synthetic dataset, shows improvements on 7 out of 8 REC datasets. Through experiments and synthetic data analysis, our findings are: (1) current MLLMs can serve as robust data generators without assistance from GPT-4V; (2) MLLMs trained with task-specific datasets can surpass GPT-4V in generating complex instruction tuning data; (3) synthetic datasets enhance performance across various multimodal benchmarks and help mitigate model hallucinations. The data, code, and models can be found at https://github.com/zhaohengyuan1/Genixer.
@@
Multimodal Large Language ModelInstruction Tuning1 Introduction
Large Language Models (LLMs) [6, 52, 2, 20] have achieved remarkable success in tackling complex natural language tasks. This progress has recently extended to Multimodal Large Language Models (MLLMs) [25, 55, 17, 40, 78, 20, 34, 10, 74, 63, 80, 3, 21]. These MLLMs demonstrate an exceptional ability to solve various multimodal problems, prompting the natural question: “Are these MLLMs capable of generating multimodal data?” This question arises due to the continued reliance on labor-intensive human labeling and the high cost of using models like GPT-4(V) for prompts. To address this gap, this paper is the first to explore the potential of current MLLMs in data generation.
Recent MLLMs [39, 17, 9, 3, 14, 21, 10, 68] demonstrate that the visual instruction data is essential for multimodal learning. Currently, the data used for model training primarily comes from two methods: (1) reformulating current vision-language datasets into instruction-following format and (2) prompting GPT-4 to create the visual instruction datasets. For the first way, models like InstructBLIP [17] trained on traditional VL tasks, such as visual question answering (VQA) and image captioning, demonstrate effectiveness on downstream tasks. However, the datasets of these tasks suffer from a limitation in image diversity, as most of them originate from the COCO dataset [38], potentially restricting the model generalization ability. For the other way, LLaVA [39], Shikra [10], and ShareGPT4V [11] propose prompting GPT-4 to produce the additional visual instruction tuning data. The primary limitations of prompting GPT-4 are twofold: (1) high financial costs for large-scale dataset creation. (2) inferior performance on some complex tasks such as Referring Expression Comprehension (REC) as illustrated in Fig. 1. Considering these issues, we propose exploring an alternative approach to generating visual instruction tuning data by training an MLLM exclusively for data generation. The benefits of this approach are twofold: (1) it eliminates additional financial costs for data generation, and (2) it provides flexibility in producing high-quality visual instruction data for arbitrary unlabeled images.
In this study, addressing the problem of visual instruction tuning data generation we consider two modes: task-specific visual data generation and task-agnostic data generation. Task-specific data generation is used to build synthetic datasets tailored to explicit task types, while task-agnostic data generation converts unlabeled images into various task types. To achieve this goal, we introduce a new pipeline Genixer, as illustrated in Fig. 2. This pipeline consists of four key steps: (i) instruction data collection, (ii) instruction template design, (iii) empowering MLLM, and (iv) data generation and filtering. in the first step, we focus on collecting prevalent and representative vision-language (VL) tasks as sources for data generation. The second step involves meticulously designing two-level instruction templates to enable controlled data generation for both task-specific and task-agnostic. In the third step, we select two representative MLLMs, LLaVA1.5, and Shikra, to cover two main VL task groups data generation: generic tasks and grounding tasks. Both MLLMs are trained to generate instruction data based on their possessed multimodal understanding. In the fourth step, to ensure the quality of the synthetic data, we propose removing incorrect data samples using two newly introduced automatic data filtering pipelines.
Using the proposed Genixer, we build two visual instruction tuning datasets: Genixer-915K for VQA tasks and Genixer-350K for REC tasks. Our experimental results show that training with 915K VQA-like data improves LLaVA1.5 [39] on 10 out of 12 multimodal datasets and benchmarks, such as 3.8% improvement on Vizwiz [23], 2.7% on SicenceQA [46], 37.7% on MME. Additionally, learning with 350K REC-like synthetic data, Shikra shows consistent improvements on 7 out of 8 REC datasets. Beyond quantitative evaluation, human study and dataset analysis, as detailed in Sec. 4.1 and 4.2, demonstrate the diversity and effectiveness of generated instruction tuning data. Visualizing examples of different task types in Fig. 3, it is evident that our Genixer successfully produces long and meaningful answers on some challenging tasks such as Multi-turn Dialogue, PointQA, and Referential Dialogue.
Overall, our approach yields three main findings: (1) MLLMs trained using our pipeline can produce visual instruction data of comparable quality without the assistance of GPT-4V. (2) MLLMs trained with our pipeline excel at generating more complex instruction tuning data compared to GPT-4V. (3) Synthetic datasets significantly enhance MLLMs’ performance on various multimodal benchmarks and help mitigate hallucinations in the models. In summary, our contributions are as follows:
We present a holistic data generation pipeline, Genixer, capable of generating diverse visual instruction tuning data from unlabeled images.
We contribute two open-source data generator models GenixerL and GenixerS for advancing data creation in multimodal domains.
We contribute two high-quality multimodal datasets, Genixer-915K and Genixer-350K, for enhancing other MLLMs in multiple benchmarks.
@@ 2 Related Work
Multimodal Large Language Models. Large Language Models (LLMs) showcased the remarkable complicated reasoning abilities. Some well-known open-souced LLMs include FlanT5 [16], OPT [79], LLaMA [65], Vicuna [15] and LLaMA-2 [66] show exceptional reasoning ability of solving math, codes problems. By leveraging these LLMs, a surge of multimodal modes [36, 40, 17, 3, 10, 68, 80, 9, 8, 47, 69, 25, 55, 51, 74, 20, 43, 24, 32, 7, 44] are proposed to integrate the visual information for diverse multimodal reasoning tasks such as image captioning [12, 1, 76] and visual question answering [22, 50, 26, 23]. LLaVA [40] is a pioneering approach that adopts a single linear layer to project the visual features extracted by CLIP [57] encoder to the input of LLM. Different from LLaVA, InstructBLIP [17] employs an instruction-aware feature extractor and obtains advanced performance on various tasks building upon the pretrained BLIP2 [36]. Besides focusing on traditional multimodal tasks, some studies [75, 10, 8, 72, 58, 59] focus on solving the grounding tasks with the power of LLMs. Shikra [10] and Ferret [8] pay attention to curating the visual instruction data, and PVIT [8] employs the region-based vision encoder. Except for these MLLMs, CogVLM [68], Qwen-VL [3], and Kosmos-2 [55] explore adopting a billion-scale pertaining data corpus to enhance the model generalization ability and robustness.
Multimodal Instruction Data. High-quality visual instruction data is crucial for training an MLLM. Two main setups are as follows: 1) Transforming the current vision-language datasets into the instruction tuning format, e.g., InstructBLIP. Such choice is limited by the diversity of image sources. 2) Some approaches LLaVA, VisionLLM[69], and Shikra, resort to prompting the GPT-4 [52] language model to generate corresponding instruction data. This way requires the image datasets to have enough captions or region-based annotations (e.g., bounding boxes), which heavily restricts the data scale. Additionally, prompting commercial LLMs incurs high costs, and even GPT-4V [53] may not address the data generation effectively on some specific tasks, as illustrated in Fig. 1. Recently, some works [70, 71] in natural language processing domains propose to generate text instruction tuning datasets for performance improvements. In the multimodal domain, some works have also been proposed to address image caption data generation [11] and text-centric instruction tuning data generation [64]. However, both approaches rely on prompting GPT-4V or Gemini Pro to build their synthetic datasets. Unlike these works, we introduce Genixer, an innovative pipeline that explores the capabilities of MLLMs to generate high-quality visual instruction data without assistance from current commercial LLMs. Our approach is the first to demonstrate the effectiveness of training current MLLMs to generate task-specific visual instructing tuning data and improve the MLLMs on several multimodal datasets and benchmarks.
3 Genixer: An Automatic Visual Instruction Tuning Data Generation Pipeline
@@
| Generic | Common VQA | VQAv2, GQA, Counting110K†, POPE‡ | 583K |
| Adv VQA | POPE‡ | 20K | |
@@
Though current MLLMs [39, 3, 68] show exceptional capability of handling various multimodal tasks, rare works concentrate on visual instruction tuning data generation. To this end, We propose Genixer, as illustrated in Fig. 2, which is a novel pipeline that contains four key steps, including 1) instruction data collection, 2) instruction template design, 3) empowering MLLMs, and 4) data generation and filtering. In the following, we will elaborate on these four key steps.
@@ 3.1 Instruction Data Collection
In accordance with the prevalence and practical relevance of real-world VL tasks, we carefully select nine representative tasks as listed in Tab. 1, including Common Visual Question Answering (Common VQA), Adversarial-based VQA (Adv VQA), Multi-choice VQA (MC VQA), Multi-turn Dialogue (MD), Referring Expression Comprehension (REC), Referring Expression Generation (REG), PointQA, QCBoxA and Referential Dialogue (RD). We divide these tasks into two main categories, generic tasks and grounding tasks.
@@ 3.2 Instruction Template Design
In an automatic data generation context, where image content is agnostic, preemptively determining the specific task type becomes particularly daunting, especially when it involves large-scale data creation purposes. Hence, we consider two key modes for visual instruction data generation: 1) task-agnostic data generation and 2) task-specific data generation.
Two-level Instructions. We propose a two-level instruction template for controlling the task type of generated visual instruction tuning data. The instruction template is as follows:
@@
Regarding Generic Instruction, it allows the model to generate arbitrary types of instruction tuning data referred to as mode 1. During training, we randomly select one of 58 handwritten instructions each time. For instance, “Please provide a clear and direct question and answer after examining the image”. Then, for Specific Instruction, it is designed to determine the specific task type, such as “This is a Common VQA task”, allowing us to control the task type referred to as mode 2.
Controlling Constant. During the training phase, we set a constant for controlling the ratio of training samples that are exclusive with Generic Instruction. Consequently, in the inference phase, the model is able to switch the mode by adding specific instructions or not. For example, as illustrated in Fig. 4, the model is capable of generating various types of data in the absence of specific instructions. Simultaneously, it can produce specific outputs when guided by a detailed prompt, like “This is an MC VQA task”.
@@ 3.3 Empowering Current MLLMs
To train an MLLM with the ability of data generation, we leverage the two remarkable MLLMs, LLaVA1.5 and Shikra, as the backbone models for generic and grounding task generation. Consequently, we obtain two data generators, GenixerL and GenixerS.
Overall Framework of GenixerL and GenixerS. For brevity, we denote the MLLM model as , the generic and specific instructions as and . Then, given an image , training objective is to make to generate question and corresponding answer :
@@
where is the whole sentence compose and , and is current prediction token. is the length of the model response sequence. denotes the total trainable parameters in (e.g., the parameters of projector and LLM with LLaVA1.5 backbone).
Training of GenixerL. GenixerL trains the pretrained LLaVA1.5 [39] for four kinds of generic tasks, including Common VQA, Adv VQA, MC VQA, and MD. As summarized in Tab. 1, we only sample a subset of original datasets from these task types for training efficiency. The controlling constant is set to 0.2, 0.2, 0.5, and 0.2, respectively. The different ratios are because of the different data sizes of these task types. These values are chosen manually to keep the training data balanced. Finally, we use AdamW [42] optimizer with a learning rate of and a batch size of 128 for one epoch training takes about 14 hours.
Training of GenixerS. We utilize the Shikra as the backbone model to train an MLLM with the ability to generate grounding task data. As shown in Tab. 1, the dataset size for RD is relatively small. To counteract potential biases in the generation, GenixerS that finetunes a SoTA model Shikra [10] on region-based tasks adopts two-phased training. The first phase focuses on the REC and REG data generation. In the second phase, GenixerS adds PointQA, QCBoxA, and RD data while deliberately reducing REC and REG data for data trade-off. To train the GenixerS, we set the equal to 0.5 for each task type. We utilize the AdamW optimizer, applying learning rates of for the first phase and for the second phase. The batch sizes are set to 128 and 64, respectively.